A tank can be fully filled with water using a pipe that fills 20 litres in a minute. A bigger pipe that can fill 25 litres in a minute will take one minute less to fill the same tank. How many minutes does the smaller pipe take to fill the tank?
Found 2 solutions by mananth, jrfrunner:
Answer by mananth(16946)

Time small pipe - time big pipe = 1 minute
...
Let x be the capacity of tank in liters
x/20-x/25= 1
(5x-4x)/100 =1
x= 100 liters
...
m.ananth@hotmail.ca

Answer by jrfrunner(365)

let t=time for smaller pipe to fill tank
let t-1 =time for larger pipe to fill tank
let c=tank capacity
==
Given
small pipe rate = 20 liters per minute
large pipe rate=25 liters per minute
==
use formula: tank capacity=rate * time
--
small pipe; tank capacity c=20*t
large pipe: tank capacity c=25*(t-1)
--
since tank capacity is the same for both pipes
20*t=25*(t-1)
20*t=25*t-25
25=5*t
5=t
--
answer: it takes the small pipe 5 minutes to fill the tank